# OUJ-FTC-21 OCHA-PP-384

### v3

@@ Abstract
We apply Nambu non-equilibrium thermodynamics (NNET)—a dynamics with multiple Hamiltonians coupled to entropy-induced dissipation—to paradigmatic far-from-equilibrium systems. Concretely, we construct NNET realizations for the Belousov–Zhabotinsky (BZ) reaction (oscillatory), the Hindmarsh–Rose neuron model (spiking), and the Lorenz and Chen systems (chaotic), and analyze their dynamical and thermodynamic signatures. Across all cases the velocity field cleanly decomposes into a reversible Nambu part and an irreversible entropy-gradient part, anchored by a model-independent quasi-conserved quantity. This construction reproduces cycles, spikes, and strange-attractor behavior and clarifies transitions among steady, periodic, and chaotic regimes via cross-model diagnostics. These results demonstrate that NNET provides a unified, quantitatively consistent framework for oscillatory, spiking, and chaotic non-equilibrium systems, offering a systematic description beyond the scope of linear-response theories such as Onsager’s relations or GENERIC.
@@ 1 Introduction
Classical non-equilibrium thermodynamics and self-organization, ranging from de Groot–Mazur[1] and the Glansdorff–Prigogine theory[2] to Nicolis–Prigogine[3] and Haken’s synergetics[4], address irreversibility, stability, and pattern selection. The GENERIC formalism[5, 6, 7, 8] further unifies reversible and irreversible couplings within a bracket structure. In contrast, our NNET makes the reversible–irreversible [^1] In this work, “reversible” does not mean microscopic time-reversal symmetry in the strict sense. Rather, it refers to dynamical reversibility: when the dissipative term is removed, the resulting (dissipationless) flow is invertible so that the admissible backward trajectory is uniquely determined by the forward one. Positivity constraints on variables do not by themselves invalidate this notion; for instance, dissipationless Lotka–Volterra dynamics remains dynamically reversible even under a positive-domain restriction. This is analogous to particle motion described in polar coordinates, which remains reversible despite coordinate or domain constraints. split explicit via Nambu brackets plus an entropy gradient. For rhythms near stable limit cycles, phase reduction originated with Winfree[9] and Kuramoto[10] and has since been systematized[11, 12]. Extensions that incorporate amplitude degrees of freedom via isostables and Koopman theory[13, 14] enlarge the regime beyond weak forcing; for large oscillator ensembles, the Ott–Antonsen ansatz[15] provides low-dimensional closures. Yet these approaches remain essentially local and do not furnish a thermodynamic partition of reversible versus dissipative contributions. Spatially extended pattern formation is grounded in reaction–diffusion theory from Turing[16] to modern reviews[17, 18]. For chemical oscillations, the Belousov-Zhabotinsky (BZ) reaction is canonically treated by the Oregonator mechanism and comprehensive monographs[19, 20, 21]. On the life-science side, spiking/bursting spans biophysically grounded models[22], geometric classifications[23], and modern syntheses[24]. These literatures motivate a unified far-from-equilibrium decomposition, which we supply via NNET.
@@
To address such systems, we have proposed Nambu Non-equilibrium Thermodynamics (NNET), a framework based on Nambu brackets that couples multiple Hamiltonians with entropy-induced dissipation. In this formulation, the velocity field separates cleanly into a reversible Nambu component and an irreversible entropy-gradient component, organized by a model-independent quasi-conserved quantity (i.e., a slowly varying variable that behaves as conserved on the relevant timescale, such as catalyst concentration in chemical oscillations or the bursting variable in neuron models; see [25] for a formal definition).
@@
Building on the original study “Fluctuating Non-linear Non-equilibrium System in Terms of Nambu Thermodynamics”[26], which integrates reversible Nambu dynamics[27] with entropy-driven dissipation, this paper further develops the framework by broadening and strengthening its applications to concrete models. As contributions of this paper, by analyzing the time evolution of representative models—the Belousov–Zhabotinsky reaction[29, 30] (oscillations), the Hindmarsh–Rose neuron[31] (spiking), and the Lorenz[32] and Chen[33] systems (chaos)—within the Nambu formalism, we demonstrate how Hamiltonian-derived conservative dynamics and entropy-driven dissipation coexist and interact. This establishes that such diverse non-equilibrium behaviors can be understood at least in the paradigmatic cases studied here as manifestations of Nambu Non-equilibrium Thermodynamics (NNET), providing a unifying description beyond the scope of linear-response theories such as those of Onsager[34] or GENERIC[6, 7, 8], in particular by accommodating regimes where entropy can locally decrease as part of the reversible–irreversible interplay.
@@
The remainder of this paper is organized as follows: In Section 2, we describe the construction of Hamiltonians and entropy within the Nambu formalism, restricting throughout to the case $N=3$ for concreteness. The general case of arbitrary $N$ has already been discussed in [28]. Section 3 presents practical applications and numerical analyses for the BZ reaction, the Hindmarsh–Rose[31] model, the Lorenz system[32], and the Chen system[33]. Finally, in Section 4, we conclude with a summary and discussion of future directions. Appendices collect supplementary materials: NNET foundations in summary (for the present 3D applications), a minimal toy model reproducing cycle–spike transitions, a near-harmonic approximation for the BZ limit cycle, and spike onset geometry analytical details.
@@ 2.1 General evolution law of NNET
In this section, we provide an overview of the Nambu Non-equilibrium Thermodynamics (NNET) framework. The Nambu Non-equilibrium Thermodynamics (NNET) framework couples two essential flows: a reversible, volume-preserving flow generated by Nambu dynamics with multiple Hamiltonians, and an irreversible flow driven by the gradient of an entropy-like potential [^2] In Onsager’s near-equilibrium theory, this $S$ coincides with the thermodynamic entropy, but in general it can differ.($S$). This dual structure cleanly partitions the dynamics into conservative and dissipative contributions and furnishes a thermodynamic interpretation of the latter. In what follows, we first present the general evolution law (here restricted to ($N=3$), then give constructive procedures for identifying the Hamiltonians and the entropy [^3] This generalized $S$ can also be interpreted as a Lyapunov function or a dissipation potential.($S$) for a given system, and finally illustrate the recipe on concrete models. A detailed axiomatic foundation is provided in [25].
We begin by recalling the basic objects of the NNET formulation and then state the general evolution law. In NNET, a system is described by an $N$-dimensional thermodynamic state vector $x=(x^{1},\dots,x^{N})$. The dynamics is decomposed into (i) a reversible, volume-preserving flow generated by multiple Hamiltonians $H_{1},\dots,H_{N-1}$ via a Nambu bracket, and (ii) an irreversible flow driven by the gradient of an entropy-like potential $S$. Accordingly, the time evolution of the state variables is given by
@@
$\{A_{1},A_{2},\cdots,A_{N}\}_{\mathrm{NB}}\equiv\frac{\partial(A_{1},A_{2},\cdots,A_{N})}{\partial(x_{1},x_{2},\cdots,x_{N})},$ (2)
governs the conservative part of the dynamics, while the gradient term represents the dissipative contribution. Here and below, $\partial^{\,i}S$ denotes the $i$-th component of $\nabla S$.
@@ 2.2 Constructive procedure for Hamiltonians and entropy
$\begin{cases}\dot{x}^{1}=2\partial_{2}H_{1}+\partial_{1}S,\\ \dot{x}^{2}=-2\partial_{1}H_{1}+\partial_{2}S,\\ \dot{x}^{3}=\partial_{3}S.\end{cases}$ (9)
@@
$2\partial_{2}^{2}H_{1}+2\partial_{1}^{2}H_{1}$ $=$ $\partial_{2}\dot{x}^{1}-\partial_{1}\dot{x}^{2},$ (10)
$2\partial_{3}\partial_{2}H_{1}$ $=$ $\partial_{3}\dot{x}^{1}-\partial_{1}\dot{x}^{3},$ (11)
$-2\partial_{3}\partial_{1}H_{1}$ $=$ $\partial_{3}\dot{x}^{2}-\partial_{2}\dot{x}^{3}.$ (12)
@@
$\left(2\epsilon_{kj3}\frac{\partial^{2}H_{1}}{\partial x^{i}\partial x^{j}}-2\epsilon_{ij3}\frac{\partial^{2}H_{1}}{\partial x^{k}\partial x^{j}}+\partial_{k}\dot{x}_{i}-\partial_{i}\dot{x}_{k}\right)dx^{k}\wedge dx^{i}=0$ (17)
@@ 3.1 BZ reaction
As a representative periodic system, we reconstruct the Belousov–Zhabotinsky (BZ) reaction within the NNET framework and discuss the time dependence of the Hamiltonians ($H_{i}$) and the entropy ($S$). The BZ reaction[29, 30] is a typical case of a time-oscillating chemical reaction. The BZ reaction is a prototypical example of a non-equilibrium chemical oscillation, in which the concentrations of intermediates exhibit periodic temporal changes accompanied by striking color oscillations. As emphasized in the pioneering works of Prigogine and coworkers, the BZ reaction has long been regarded as an archetype of dissipative structures that form far from equilibrium, providing key insights into the emergence of temporal order in chemical systems.
@@
Then, the equations of the time evolution of the BZ model become the equations of motion for three coordinates or three concentrations ($X$, $Y$, and $Z$), given by
@@
These chemical reactions are far from equilibrium and are not obtained from the Onsager’s variational principle.
@@
$\frac{\partial^{2}H_{1}}{\partial X^{2}}+\frac{\partial^{2}H_{1}}{\partial Y^{2}}$ $=$ $k_{1}A-k_{2}(X-Y),$ (27)
@@
$\frac{\partial f(X,Y)}{\partial X}+\frac{\partial g(X,Y)}{\partial X}$ $=$ $k_{1}A-k_{2}(X-Y).$
@@
$H_{1}=hk_{5}BZX-\frac{1}{2}k_{1}AY^{2}+\frac{k_{2}}{6}Y^{3}+2k_{3}AYZ+\frac{1}{6}k_{2}X^{3}.$ (31)
The time evolution of the non-Hamiltonian part is necessarily in the form of an entropy gradient. To derive the entropy $S$, we use the decomposition of the velocity field $v$ in Eq. (6) into an incompressible Nambu part and a gradient part, namely., $v=v^{(H)}+\nabla S$. Concretely, after determining $H_{1}$, we define the Nambu (Hamiltonian) contribution to the velocities by
@@
and the residual by $v^{(S)}:=\dot{x}-v^{(H)}$. The consistency relation Eq. (15) ensures that this residual is an exact gradient, $v^{(S)}=\nabla S$, which is also equivalent to $\nabla\cdot v=\nabla^{2}S$ in Eq. (5). Integrating $\partial_{i}S=v^{(S)}_{i}$ (up to an additive constant) yields the explicit expression. This construction follows [25, 28].
@@
$\partial_{t}^{(H)}Z$ $=$ $-\{H_{1},H_{2},Z\}_{\mathrm{NB}}=0,$ (38)
while the time evolution induced by the entropy is given by
@@
$\partial_{t}^{(S)}S$ $=$ $+\{S,X,Y\}_{\mathrm{NB}}^{2}+\{S,Y,Z\}_{\mathrm{NB}}^{2}+\{S,Z,X\}_{\mathrm{NB}}^{2}.$ (44)
The entropic part of the entropy time evolution is always positive, reflecting the essence of entropy production. In contrast, the Hamiltonian part may become negative, as it corresponds to reversible dynamics. Such negative contributions can arise, for example, in oscillatory reactions like the BZ reaction or in non-equilibrium steady states, where reversible energy or matter redistribution leads to a temporary decrease in entropy.
@@
$\{H_{1},H_{2},S\}_{\mathrm{NB}}=\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\partial_{t}^{(H)}X\partial_{t}^{(S)}X\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\partial_{t}^{(H)}Y\partial_{t}^{(S)}Y\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\partial_{t}^{(H)}Z\partial_{t}^{(S)}Z.$ (45)
@@
$\dot{S}$ $=$ $-\partial_{t}^{(H)}X\partial_{t}^{(S)}X-\partial_{t}^{(H)}Y\partial_{t}^{(S)}Y-\partial_{t}^{(H)}Z\partial_{t}^{(S)}Z\$(46)
@@
$\{H_{1},X,Y\}_{\mathrm{NB}}=-k_{5}hBX+2k_{3}AY,$ (50)
$\{H_{1},Y,Z\}_{\mathrm{NB}}=\frac{k_{2}X^{2}}{2}-k_{5}hBZ,$ (51)
@@
Note that oscillations do not occur in all parameter regions. In particular, when the parameter relation $hk_{5}\sim k_{1}A\sim k_{2}\sim k_{4}X\sim k_{5}B\ll k_{3}A$ holds, the BZ reaction exhibits oscillatory behavior. As discussed in Appendix C, under this condition the system can be approximated by a set of harmonic oscillator Hamiltonians, which provides a clear theoretical basis for the emergence of oscillations. Approximating the Hamiltonian and entropy in the case, we get
@@
This represents a system of harmonic oscillators as discussed in Appendix C. Thus, under the appropriate variable transformations, the system is approximately described as a system of harmonic oscillator Hamiltonians with no entropy. Thus, from the given Hamiltonian and entropy, we can analyze whether oscillatory phenomena occur [^5] When the condition $dS=dH_{1}\wedge d\Delta H_{2}+d\Delta H_{1}\wedge dH_{2}+d\Delta H_{1}\wedge d\Delta H_{2}$ is satisfied, the redefinition $H_{i}\rightarrow H_{i}+\Delta H_{i}$ leads to the elimination of the entropy $S$, leaving only the Hamiltonians to describe the system. This removal of $S$ is crucial in achieving a quasi-conserved system of Hamiltonians. In general, the interaction between the set of Hamiltonians and entropy plays a significant role in determining the system’s behavior..

### v4

# arXiv:2509.12641v4[cond-mat.stat-mech] 28 Feb 2026

@@ Abstract
We apply Nambu non-equilibrium thermodynamics (NNET)—a dynamics with multiple Hamiltonians coupled to entropy-induced dissipation—to paradigmatic far-from-equilibrium systems. Concretely, we construct NNET realizations for the Belousov–Zhabotinsky (BZ) reaction (oscillatory), the Hindmarsh–Rose neuron model (spiking), and the Lorenz and Chen systems (chaotic), and analyze their dynamical and thermodynamic signatures. Across all cases the velocity field cleanly decomposes into a non-dissipative Nambu part and an dissipative entropy-gradient part, anchored by a model-independent quasi-conserved quantity. This construction reproduces cycles, spikes, and strange-attractor behavior and clarifies transitions among steady, periodic, and chaotic regimes via cross-model diagnostics. These results demonstrate that NNET provides a unified, quantitatively consistent framework for oscillatory, spiking, and chaotic non-equilibrium systems, offering a systematic description beyond the scope of linear-response theories such as Onsager’s relations or GENERIC.
@@ 1 Introduction
Classical non-equilibrium thermodynamics and self-organization, ranging from de Groot–Mazur[1] and the Glansdorff–Prigogine theory[2] to Nicolis–Prigogine[3] and Haken’s synergetics[4], address dissipative, stability, and pattern selection. The GENERIC formalism[5, 6, 7, 8] further unifies non-dissipative and dissipative couplings within a bracket structure. In contrast, our NNET makes the non-dissipative and dissipative  [^1] In this work, “non-dissipative” does not imply microscopic time-reversal symmetry. It refers to the structural property that, after removing the entropy-gradient term, the remaining Nambu flow is volume-preserving (incompressible) and generates a well-posed deterministic flow (unique forward and backward trajectories) within the admissible state domain. split explicit via Nambu brackets plus an entropy gradient. For rhythms near stable limit cycles, phase reduction originated with Winfree[9] and Kuramoto[10] and has since been systematized[11, 12]. Extensions that incorporate amplitude degrees of freedom via isostables and Koopman theory[13, 14] enlarge the regime beyond weak forcing; for large oscillator ensembles, the Ott–Antonsen ansatz[15] provides low-dimensional closures. Yet these approaches remain essentially local and do not furnish a thermodynamic partition of non-dissipative versus dissipative contributions. Spatially extended pattern formation is grounded in reaction–diffusion theory from Turing[16] to modern reviews[17, 18]. For chemical oscillations, the Belousov-Zhabotinsky (BZ) reaction is canonically treated by the Oregonator mechanism and comprehensive monographs[19, 20, 21]. On the life-science side, spiking/bursting spans biophysically grounded models[22], geometric classifications[23], and modern syntheses[24]. These literatures motivate a unified far-from-equilibrium decomposition, which we supply via NNET.
@@
To address such systems, we have proposed Nambu Non-equilibrium Thermodynamics (NNET), a framework based on Nambu brackets that couples multiple Hamiltonians with entropy-induced dissipation. In this formulation, the velocity field separates cleanly into a non-dissipative Nambu component and an dissipative entropy-gradient component, organized by a model-independent quasi-conserved quantity (i.e., a slowly varying variable that behaves as conserved on the relevant timescale, such as catalyst concentration in chemical oscillations or the bursting variable in neuron models; see [25] for a formal definition).
@@
Building on the original study “Fluctuating Non-linear Non-equilibrium System in Terms of Nambu Thermodynamics”[26], which integrates non-dissipative Nambu dynamics[27] with entropy-driven dissipation, this paper further develops the framework by broadening and strengthening its applications to concrete models. As contributions of this paper, by analyzing the time evolution of representative models—the Belousov–Zhabotinsky reaction[29, 30] (oscillations), the Hindmarsh–Rose neuron[31] (spiking), and the Lorenz[32] and Chen[33] systems (chaos)—within the Nambu formalism, we demonstrate how Hamiltonian-derived non-dissipative dynamics and entropy-driven dissipation coexist and interact. This establishes that such diverse non-equilibrium behaviors can be understood at least in the paradigmatic cases studied here as manifestations of Nambu Non-equilibrium Thermodynamics (NNET), providing a unifying description beyond the scope of linear-response theories such as those of Onsager[34] or GENERIC[6, 7, 8], in particular by accommodating regimes where entropy can locally decrease as part of the non-dissipative and dissipative interplay.
@@
The remainder of this paper is organized as follows: Section 2 includes a compact summary of the NNET formulation for the present 3D applications. The general case of arbitrary $N$ has already been discussed in [28]. Section 3 presents practical applications and numerical analyses for the BZ reaction, the Hindmarsh–Rose[31] model, the Lorenz system[32], and the Chen system[33]. Finally, in Section 4, we conclude with a summary and discussion of future directions. The appendices collect supplementary materials: a minimal toy model reproducing cycle–spike transitions, a near-harmonic approximation for the BZ limit cycle, and analytical details on spike onset geometry.
@@ 2.1 General evolution law of NNET
In this section, we briefly review the Nambu Non-equilibrium Thermodynamics (NNET) framework. NNET couples two essential flows: a non-dissipative, volume-preserving flow generated by Nambu dynamics with multiple Hamiltonians, and a dissipative flow driven by the gradient of an entropy-like potential [^2] In Onsager’s near-equilibrium theory, this $S$ coincides with the thermodynamic entropy, but in general it can differ.($S$). This dual structure cleanly partitions the dynamics into conservative and dissipative contributions and furnishes a thermodynamic interpretation of the latter. In what follows, we first present the general evolution law (here restricted to $N=3$), then give constructive procedures for identifying the Hamiltonians and the entropy [^3] This generalized $S$ can also be interpreted as a Lyapunov function or a dissipation potential.($S$) for a given system, and finally illustrate the recipe on concrete models. A detailed axiomatic foundation is provided in [25].
We begin by recalling the basic objects of the NNET formulation and then state the general evolution law. In NNET, a system is described by an $N$-dimensional thermodynamic state vector $x=(x^{1},\dots,x^{N})$. The dynamics is decomposed into (i) a non-dissipative, volume-preserving flow generated by multiple Hamiltonians $H_{1},\dots,H_{N-1}$ via a Nambu bracket, and (ii) an dissipative flow driven by the gradient of an entropy-like potential $S$. Accordingly, the time evolution of the state variables is given by
@@
$\{A_{1},A_{2},\cdots,A_{N}\}_{\mathrm{NB}}\equiv\frac{\partial(A_{1},A_{2},\cdots,A_{N})}{\partial(x^{1},x^{2},\cdots,x^{N})},$ (2)
governs the non-dissipative part of the dynamics, while the gradient term represents the dissipative contribution. Here and below, $\partial_{i}S$ denotes the $i$-th component of $\nabla S$.
@@
2.2 NNET in summary (for the present 3D applications)
We consider an $n$-dimensional state vector $\bm{x}=(x^{1},\dots,x^{n})$ defined on a regular region of the state space. Following the axiomatic formulation of NNET [25], the time evolution is decomposed into a non-dissipative part and a dissipative part,
$\dot{x}^{i}=v_{\mathrm{non\text{-}diss}}^{i}(\bm{x})+v_{\mathrm{diss}}^{i}(\bm{x}),\qquad(i=1,\dots,n).$ (6)
The non-dissipative part is generated by a Nambu bracket with multiple Hamiltonians, while the dissipative part is a gradient flow driven by an entropy-like potential $S$ with a positive-definite coefficient (metric) $g^{ij}$.
In the three-dimensional systems considered in this paper ($n=3$), the Nambu bracket is defined as the Jacobian
$\{F,G,H\}_{\mathrm{NB}}:=\frac{\partial(F,G,H)}{\partial(x^{1},x^{2},x^{3})}=\nabla F\cdot(\nabla G\times\nabla H).$ (7)
The NNET evolution law used in the main text reads, in component form,
$\dot{x}^{i}=-\{H_{1},H_{2},x^{i}\}_{\mathrm{NB}}+\sum_{j=1}^{3}g^{ij}\,\partial_{j}S,\qquad(i=1,2,3).$ (8)
In the present applications, we effectively work in a Euclidean metric, so that the dissipative term reduces to a gradient flow; the notation with $g^{ij}$ is included only to indicate the more general form.
Here “non-dissipative” is used in the structural sense: the Nambu part generates a volume-preserving (incompressible) flow and conserves the Hamiltonians in the absence of dissipation. It does not necessarily imply strict time-reversal invariance under all physical constraints.
The entropy rate is correspondingly decomposed as
$\dot{S}=-\{H_{1},H_{2},S\}_{\mathrm{NB}}+\sum_{i,j=1}^{3}(\partial_{i}S)\,g^{ij}\,(\partial_{j}S).$ (9)
If $g^{ij}$ is positive definite, the second term is non-negative and represents entropy production. The first term is not sign-definite and corresponds to the Hamiltonian (non-dissipative) contribution to the entropy rate. Therefore, in open or far-from-equilibrium situations, $\dot{S}$ need not be strictly positive.
@@ 2.3 Constructive procedure for Hamiltonians and entropy
Gauge-type non-uniqueness.
The representation of a given velocity field in terms of $(H_{1},H_{2},S)$ is not unique. If a pair $(\delta H_{1},\delta S)$ satisfies
$\{\delta H_{1},H_{2},x^{i}\}_{\mathrm{NB}}+\partial_{i}\delta S=0\qquad(i=1,2,3),$ (12)
then the transformation
$H_{1}\rightarrow H_{1}+\delta H_{1},\qquad S\rightarrow S+\delta S$ (13)
leaves the velocity field invariant. This non-uniqueness motivates the use of explicit selection principles.
Selection principles.
The gauge-type non-uniqueness described above implies that additional selection principles are required to identify a representation that is physically meaningful and useful for applications. In this work, we adopt the following criteria:
  1. 1.
    Near- vs. far-from-equilibrium regime. When a near-equilibrium limit is relevant, we impose consistency with the equilibrium structure in that limit (e.g., agreement with standard thermodynamic interpretations and Onsager-type forms when applicable).
  2. 2.
    Conserved or quasi-conserved quantities. We identify exact conserved quantities, or quantities that are approximately conserved on the time scale of interest (pseudo-conserved variables), and use them as natural candidates for one of the Hamiltonians. In far-from-equilibrium systems with attractors, such quasi-invariants can often be suggested by the slow or recurrent structure of trajectories (e.g., approximate constants of motion along the attractor), which motivates the choice of a pseudo-conserved variable in the Nambu sector.
  3. 3.
    Coupling to reservoirs and the construction of $S$. The dissipative potential $S$ is chosen in accordance with the type of reservoir (thermal, chemical, mechanical, etc.) to which the system is coupled, so that the gradient part encodes the relevant dissipation mechanism.
Once a conserved or pseudo-conserved quantity is selected as one Hamiltonian, the remaining Hamiltonian structure is then determined constructively through the Helmholtz decomposition together with the Nambu representation (via Darboux’s theorem).
As a simple illustration, consider a chemically open triangular reaction network $A\leftrightarrow B\leftrightarrow C\leftrightarrow A$ coupled to chemical reservoirs. A natural reservoir potential is the grand potential $\Omega=F-\sum_{i\in\{A,B,C\}}\mu_{i}N_{i}$, and conserved combinations (e.g., the total population $N_{\mathrm{tot}}=N_{A}+N_{B}+N_{C}$ in a closed limit, or an approximately conserved quantity on the relevant time scale in the open setting) provide natural candidates for Hamiltonians [^4] Detailed discussion of this point can be found in [25]. Once such a conserved/pseudo-conserved quantity is fixed as one Hamiltonian, the remaining Hamiltonian is then constrained by the requirement that the Nambu (solenoidal) component reproduces the rotational part of the velocity field.
In practice, throughout this paper we adopt the simplest choice of $(H_{1},H_{2},S)$ that reproduces the solenoidal and gradient parts of the given vector field and is consistent with the relevant conserved or quasi-conserved quantities.
@@
$\begin{cases}\dot{x}^{1}=\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\partial_{2}H_{1}+\partial_{1}S,\\ \dot{x}^{2}=\color[rgb]{0,0,0}\partial_{1}H_{1}+\partial_{2}S,\\ \dot{x}^{3}=\partial_{3}S.\end{cases}$ (15)
@@
$\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\partial_{2}^{2}H_{1}\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\partial_{1}^{2}H_{1}$ $=$ $\partial_{2}\dot{x}^{1}-\partial_{1}\dot{x}^{2},$ (16)
$\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\partial_{3}\partial_{2}H_{1}$ $=$ $\partial_{3}\dot{x}^{1}-\partial_{1}\dot{x}^{3},$ (17)
$\color[rgb]{0,0,0}\partial_{3}\partial_{1}H_{1}$ $=$ $\partial_{3}\dot{x}^{2}-\partial_{2}\dot{x}^{3}.$ (18)
@@
$\left(\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\epsilon_{kj3}\frac{\partial^{2}H_{1}}{\partial x^{i}\partial x^{j}}\color[rgb]{0,0,0}+\color[rgb]{0,0,0}\epsilon_{ij3}\frac{\partial^{2}H_{1}}{\partial x^{k}\partial x^{j}}+\partial_{k}\dot{x}_{i}-\partial_{i}\dot{x}_{k}\right)dx^{k}\wedge dx^{i}=0$ (23)
@@ 3.1 BZ reaction
As a representative periodic system, we reconstruct the Belousov–Zhabotinsky (BZ) reaction within the NNET framework and discuss the time dependence of the Hamiltonians ($H_{i}$) and the entropy ($S$) The BZ reaction[29, 30] is a prototypical example of a non-equilibrium chemical oscillation, in which the concentrations of intermediates exhibit periodic temporal changes accompanied by striking color oscillations. As emphasized in the pioneering works of Prigogine and coworkers, the BZ reaction has long been regarded as an archetype of dissipative structures that form far from equilibrium, providing key insights into the emergence of temporal order in chemical systems.
@@
The time-evolution equations for the three concentrations $(X,Y,Z)$ are
@@
These chemical reactions are far from equilibrium and are not obtained from Onsager’s variational principle.
@@
$\frac{\partial^{2}H_{1}}{\partial X^{2}}+\frac{\partial^{2}H_{1}}{\partial Y^{2}}$ $=$ $\color[rgb]{0,0,0}-\color[rgb]{0,0,0}k_{1}A\color[rgb]{0,0,0}+\color[rgb]{0,0,0}k_{2}(X-Y),$ (33)
@@
$\frac{\partial f(X,Y)}{\partial X}+\frac{\partial g(X,Y)}{\partial X}$ $=$ $\color[rgb]{0,0,0}-\color[rgb]{0,0,0}k_{1}A\color[rgb]{0,0,0}+\color[rgb]{0,0,0}k_{2}(X-Y).$
@@
$H_{1}=hk_{5}BZX-\frac{1}{2}k_{1}AY^{2}\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\frac{k_{2}}{6}Y^{3}+2k_{3}AYZ+\frac{1}{6}k_{2}X^{3}.$ (37)
The time evolution of the non-Hamiltonian part is necessarily in the form of an entropy gradient. To derive the entropy $S$, we use the decomposition of the velocity field $v$ in Eq. (10) into an incompressible Nambu part and a gradient part, namely, $v=v^{(H)}+\nabla S$. Concretely, after determining $H_{1}$, we define the Nambu (Hamiltonian) contribution to the velocities by
@@
and the residual by $v^{(S)}:=\dot{\bm{x}}-v^{(H)}$. The consistency relation Eq. (21) ensures that this residual is an exact gradient, $v^{(S)}=\nabla S$, which is also equivalent to $\nabla\cdot v=\nabla^{2}S$ in Eq. (5). Integrating $\partial_{i}S=v^{(S)}_{i}$ (up to an additive constant) yields the explicit expression. This construction follows [25, 28].
@@
$\partial_{t}^{(H)}Z$ $=$ $-\{H_{1},H_{2},Z\}_{\mathrm{NB}}=0.$ (44)
In the present Euclidean setting, the dissipative contribution coincides with the gradient part and can be expressed equivalently in bracket form as below.
@@
$\partial_{t}^{(S)}S$ $=$ $\{S,X,Y\}_{\mathrm{NB}}^{2}+\{S,Y,Z\}_{\mathrm{NB}}^{2}+\{S,Z,X\}_{\mathrm{NB}}^{2}.$ (50)
The entropic part of the entropy time evolution is always positive, reflecting the essence of entropy production. In contrast, the Hamiltonian part may become negative, as it corresponds to non-dissipative dynamics. Such negative contributions can arise, for example, in oscillatory reactions like the BZ reaction or in non-equilibrium steady states, where non-dissipative energy or matter redistribution leads to a temporary decrease in entropy.
@@
$\{H_{1},H_{2},S\}_{\mathrm{NB}}=\color[rgb]{0,0,0}\color[rgb]{0,0,0}\partial_{t}^{(H)}X\partial_{t}^{(S)}X\color[rgb]{0,0,0}+\color[rgb]{0,0,0}\color[rgb]{0,0,0}\partial_{t}^{(H)}Y\partial_{t}^{(S)}Y\color[rgb]{0,0,0}+\color[rgb]{0,0,0}\color[rgb]{0,0,0}\partial_{t}^{(H)}Z\partial_{t}^{(S)}Z.$ (51)
@@
$\dot{S}$ $=$ $\color[rgb]{0,0,0}\partial_{t}^{(H)}X\partial_{t}^{(S)}X\color[rgb]{0,0,0}+\color[rgb]{0,0,0}\partial_{t}^{(H)}Y\partial_{t}^{(S)}Y\color[rgb]{0,0,0}+\color[rgb]{0,0,0}\partial_{t}^{(H)}Z\partial_{t}^{(S)}Z\$(52)
@@
$\{H_{1},X,Y\}_{\mathrm{NB}}=\color[rgb]{0,0,0}k_{5}hBX+2k_{3}AY,$ (56)
$\{H_{1},Y,Z\}_{\mathrm{NB}}=\frac{k_{2}X^{2}}{2}\color[rgb]{0,0,0}+\color[rgb]{0,0,0}k_{5}hBZ,$ (57)
@@
Note that oscillations do not occur in all parameter regions. In particular, when the parameter relation $hk_{5}\sim k_{1}A\sim k_{2}\sim k_{4}X\sim k_{5}B\ll k_{3}A$ holds, the BZ reaction exhibits oscillatory behavior. As discussed in Appendix B, under this condition the system can be approximated by a set of harmonic oscillator Hamiltonians, which provides a clear theoretical basis for the emergence of oscillations. Approximating the Hamiltonian and entropy in the case, we get
@@
This represents a system of harmonic oscillators as discussed in Appendix B. Thus, under the appropriate variable transformations, the system is approximately described as a system of harmonic oscillator Hamiltonians with no entropy. Thus, from the given Hamiltonian and entropy, we can analyze whether oscillatory phenomena occur [^6] When the condition $dS=dH_{1}\wedge d\Delta H_{2}+d\Delta H_{1}\wedge dH_{2}+d\Delta H_{1}\wedge d\Delta H_{2}$ is satisfied, the redefinition $H_{i}\rightarrow H_{i}+\Delta H_{i}$ leads to the elimination of the entropy $S$, leaving only the Hamiltonians to describe the system. This removal of $S$ is crucial in achieving a quasi-conserved system of Hamiltonians. In general, the interaction between the set of Hamiltonians and entropy plays a significant role in determining the system’s behavior..